# Bringing closure to FDR control: beating the e-Benjamini-Hochberg procedure

### v2

@@ 2 The closed eBH (${\overline{\text{eBH}}}$) procedure
$k^{\overline{\textnormal{eBH}}}$ $\coloneqq\max\ \{i\in[K]:R_{i}\subseteq\overline{\mathcal{C}}\}$ (5)
@@
We give a simple example with $K=3$ to show that the closed eBH rule can reject strictly more than the minimally adaptive eBH rule. For the set of input e-values $\mathbf{E}=(60,39,11)$, the eBH procedure only makes one rejection at $\alpha=0.05$, narrowly missing out on the second rejection. The minimally adaptive eBH procedure effectively runs eBH at level $23\alpha/$ 2, and makes two rejections. The closed eBH procedure rejects all 3 hypotheses, which can be checked by verifying that the set $\{1,2,3\}$ lies in $\overline{\mathcal{C}}$, since each e-value is larger than 20/3, all pairwise averages exceed 40/3 and the overall average exceeds 20.
@@
                   Let $S_{(m-r)}=\sum_{i=1}^{m-r}E_{(i)}$(sum of $m-r$ smallest e-values overall).
@@
To do this efficiently, we note that for each $r\in[k]$, we can fix $r$ to be the number of false discoveries. Then, to maximize $\textnormal{FDP}_{A}(R)/E_{A}$, $A$ includes $r$ true nulls in $R_{k}$ with the smallest e-values in $R_{k}$ and we only need to maximize over $m\in\{r,\dots,r+K-k\}$, the number of true nulls outside of the rejection set, where for each $m$ we pick the $m-r$ smallest e-values.
Let $A_{(k,r,m)}$ be the indices corresponding to the $r$ smallest e-values in $R_{k}$ and the $m-r$ smallest e-values outside of $R_{k}$. Then, we can say that
@@ 2.2 The compound ${\overline{\text{eBH}}}$ procedure
is a valid e-value for $\mathcal{H}_{A}$. Let the compound ${\overline{\text{eBH}}}$ procedure be the ${\overline{\text{eBH}}}$ procedure that utilizes (25) for its e-collection instead of (2) and be denoted as $\tilde{R}^{{\overline{\textnormal{eBH}}}}\coloneqq\mathcal{R}^{{\overline{\textnormal{eBH}}}}(\widetilde{\mathbf{E}})$. Let $\tilde{R}^{\textnormal{eBH}}\coloneqq\mathcal{R}^{\textnormal{eBH}}(\widetilde{\mathbf{E}})$ now be the discovery set that results from applying the eBH procedure.
@@ 4 Closed testing for general error metrics
$\overline{\mathcal{C}}(F)$ =$=\left\{R\subseteq[K]:E_{A}\geq\frac{\text{F}_{A}(R)}{\alpha}\quad\forall{A\subseteq[K]}\right\}$ (36)
=$=\left\{R\subseteq[K]:E_{A}\geq\frac{1}{\alpha}\text{ for all }A\text{ with }A\cap R\neq\emptyset\right\}.$ (37)
@@ 5 Extensions
$S_{A}(E_{A})$ =$=\begin{cases}\hat{\alpha}_{A,\hat{k}}^{-1}&\text{ if }\hat{\alpha}_{A,\hat{k}+1}>\hat{\alpha}_{A,\hat{k}}\text{ or }U>\frac{E_{A}-\hat{\alpha}^{-1}_{A,\hat{k}}}{\hat{\alpha}^{-1}_{A,\hat{k}+1}-\hat{\alpha}^{-1}_{A,\hat{k}}}\\ \hat{\alpha}_{A,\hat{k}+1}^{-1}&\text{ otherwise}\end{cases}.$ (43)
@@ 6 Numerical simulations
We also evaluate the BY procedure, the application of minimally adaptive eBH to the e-values defined in (2.1), and the $\overline{\text{BY}}$ procedure where we set $P_{i}\coloneqq 1-\Phi(X_{i})$ and $\Phi$ is the c.d.f. of the standard normal distribution. Here, we draw $(X_{1},\dots,X_{K})$ from a multivariate normal with the same component means as described above and with Toeplitz-like covariance matrix where $|\text{Cov}(X_{i},X_{j})|=\exp(-|i-j|/10)/5$ with the covariance being positive if $|i-j|$ is even and negative otherwise. Similarly, our results in Figure 2 show that $\overline{\text{BY}}$ has substantial power increase over the other BY procedures.

### v3

@@ 2 The closed eBH ($\overline{\text{eBH}}$}) procedure
$k^{\overline{\textnormal{eBH}}}$ $\coloneqq\max\ \{i\in[K]:R_{i}\in\overline{\mathcal{C}}\}$ (5)
@@
We give a simple example with $K=3$ to show that the closed eBH rule can reject strictly more than the minimally adaptive eBH rule. For the set of input e-values $\mathbf{E}=(60,39,11)$, the eBH procedure only makes one rejection at $\alpha=0.05$, narrowly missing out on the second rejection. The minimally adaptive eBH procedure effectively runs eBH at level $3\alpha/2$, and makes two rejections. The closed eBH procedure rejects all 3 hypotheses, which can be checked by verifying that the set $\{1,2,3\}$ lies in $\overline{\mathcal{C}}$, since each e-value is larger than 20/3, all pairwise averages exceed 40/3 and the overall average exceeds 20.
@@
         Let $S_{(m-r)}=\sum_{i=1}^{m-r}E_{(i)}$(sum of mrm-r smallest e-values overall).
@@
To do this efficiently, we note that for each $r\in[k]$, we can fix $r$ to be the number of false discoveries. Then, to maximize $\textnormal{FDP}_{A}(R)/E_{A}$, $A$ includes $r$ true nulls in $R_{k}$ with the smallest e-values in $R_{k}$ and we only need to maximize over $m\in\{r,\dots,r+K-k\}$, the number of true nulls outside of the rejection set, where for each $m$ we pick the mrm-r smallest e-values.
Let $A_{(k,r,m)}$ be the indices corresponding to the $r$ smallest e-values in $R_{k}$ and the mrm-r smallest e-values outside of $R_{k}$. Then, we can say that
@@ 2.2 The compound $\overline{\text{eBH}}$} procedure
[…]
@@ 4 Closed testing for general error metrics
$\overline{\mathcal{C}}(F)$ $=\left\{R\subseteq[K]:E_{A}\geq\frac{\text{F}_{A}(R)}{\alpha}\quad\forall{A\subseteq[K]}\right\}$ (36)
$=\left\{R\subseteq[K]:E_{A}\geq\frac{1}{\alpha}\text{ for all }A\text{ with }A\cap R\neq\emptyset\right\}.$ (37)
@@ 5 Extensions
$S_{A}(E_{A})$ $=\begin{cases}\hat{\alpha}_{A,\hat{k}}^{-1}&\text{ if }\hat{\alpha}_{A,\hat{k}+1}>\hat{\alpha}_{A,\hat{k}}\text{ or }U>\frac{E_{A}-\hat{\alpha}^{-1}_{A,\hat{k}}}{\hat{\alpha}^{-1}_{A,\hat{k}+1}-\hat{\alpha}^{-1}_{A,\hat{k}}}\\ \hat{\alpha}_{A,\hat{k}+1}^{-1}&\text{ otherwise}\end{cases}.$ (43)
@@ 6 Numerical simulations
We also evaluate the BY procedure, the application of minimally adaptive eBH to the e-values defined in (2.1), and the $\overline{\text{BY}}$ procedure where we set $P_{i}\coloneqq 1-\Phi(X_{i})$ and $\Phi$ is the c.d.f. of the standard normal distribution. Here, we draw $(X_{1},\dots,X_{K})$ from a multivariate normal with the same component means as described above and with Toeplitz-like covariance matrix where $|\text{Cov}(X_{i},X_{j})|=\exp(-|i-j|/10)/5$ with the covariance being positive if |ij||i-j| is even and negative otherwise. Similarly, our results in Figure 2 show that $\overline{\text{BY}}$ has substantial power increase over the other BY procedures.